Molds are used in the manufacture of contact lenses so that the lens material for proper preparation and curing will be consistent and meet designated dimensional criteria. Molds were fabricated and a critical dimension measured for a sample of 15 randomly selected molds. (Data have been doubly coded to ensure propriety.) Suppose the standard
deviation is known to be ? = 30.6

Find the 95% confidence interval for µ.

140 130 195 110 150 150 130 120 120 15 180 95 135 220 105
Answer #1

Sample mean calculated, x̅ = 133

Population standard deviation, σ = 30.6

Sample size, n = 15

95% Confidence interval for µ:    

At α = 0.05, two tailed critical value, z-crit = NORM.S.INV(0.05/ 2) = 1.960

Lower Bound = x̅-z-crit*σ/√n = 117.5145

Upper Bound = x̅+z-crit*σ/√n =148.4855